# The effect of priors on Learning with Restricted Boltzmann Machines

### v1

@@ Appendix A Derivation of the RS equations
$f^{RS}(\Lambda,\hat{\Lambda},\Lambda_{\tau},\hat{\Lambda}_{\tau})=f^{RS}_{0}+nf^{RS}_{1}+\mathcal{O}(n^{2})\,,$
$f^{RS}_{0}$ $=\frac{\alpha\hat{\beta}}{2}d_{\hat{\tau}}d_{\hat{\xi}}+\hat{\beta}\operatorname{\mathrm{Extr}}_{\bm{p}_{T}}f_{T}(\bm{p}_{T})+\alpha\log\mathbb{E}_{\hat{\tau}}e^{\hat{d}_{\hat{\tau}}(\hat{\tau})^{2}}+\log\mathbb{E}_{\hat{\xi}}e^{\hat{d}_{\hat{\xi}}(\hat{\xi})^{2}}-\alpha\hat{d}_{\hat{\tau}}d_{\hat{\tau}}-\hat{d}_{\hat{\xi}}d_{\hat{\xi}}\;,$
$f^{RS}_{1}$ $=\frac{1}{2}\hat{q}q+\alpha\frac{1}{2}\hat{q}_{\tau}q_{\tau}-\hat{m}m-\alpha\hat{m}_{\tau}m_{\tau}-\hat{d}d-\alpha\hat{d}_{\tau}d_{\tau}+\alpha^{2}\frac{\beta}{2}\Omega_{\xi}p_{\text{$\tau$}}^{2}+\frac{\beta}{2}\Omega_{\tau}p^{2}+$
$+\frac{\alpha}{\mathbb{E}_{\hat{\tau}}e^{\hat{d}_{\hat{\tau}}\hat{\tau}^{2}}}\Bigg{\langle}\mathbb{E}_{\hat{\tau}}e^{\hat{d}_{\hat{\tau}}\hat{\tau}^{2}}\mathbb{E}_{s}\log\mathbb{E}_{\tau}\exp\Bigg{\{}-(\frac{\hat{q}_{\tau}}{2}-\hat{d}_{\tau})\tau^{2}+\tau(\hat{m}_{\tau}\hat{\tau}+\sqrt{\hat{q}_{\tau}}z+\hat{p}_{\tau}s)\Bigg{\}}\Bigg{\rangle}_{z}+$
$+\frac{1}{\mathbb{E}_{\hat{\xi}}e^{\hat{d}_{\hat{\xi}}\hat{\xi}^{2}}}\Bigg{\langle}\mathbb{E}_{\hat{\xi}}e^{\hat{d}_{\hat{\xi}}\,\hat{\xi}^{2}}\mathbb{E}_{s}\log\mathbb{E}_{\xi}\exp\Bigg{\{}-(\frac{\hat{q}}{2}-\hat{d})\xi^{2}+\xi(\hat{m}\hat{\xi}+\sqrt{\hat{q}}z+\hat{p}s)\Bigg{\}}\Bigg{\rangle}_{z}+$
@@
The extremization of $f^{RS}_{0}$ involves only the teacher order parameters. The extremizers should be s.t. $\operatorname{\mathrm{Extr}}f^{RS}_{0}=0$ in order to avoid divergences when $n\to 0$. Indeed, remembering Eqs.(72), the parameters of order zero in $n$ satisfy the following
$\hat{d}_{\hat{\xi}}$ $=0\,,\quad d_{\hat{\xi}}=1\,,$ (89)
$\hat{d}_{\hat{\tau}}$ $=\frac{\hat{\beta}}{2}\,,\quad d_{\hat{\tau}}=\langle\hat{\tau}^{2}\rangle_{\hat{\tau}}\,,$
which implies $f^{RS}_{0}=0$ and ensure the student partition function is not diverging. Using Eqs.(79) and updating $f_{1}^{RS}$ with (89) we obtain the values of $\Lambda$ and $\Lambda_{\tau}$ that extremize $f_{1}^{RS}$
$\hat{p}_{\tau}$ $=\beta\alpha\Omega_{\xi}p_{\tau}\,,\quad\hat{m}_{\tau}=\sqrt{\beta\hat{\beta}}m\,,\quad\hat{q}_{\tau}=\beta q\,,\quad\hat{d}_{\tau}=\beta\frac{d}{2}\,,$ (90)
$\hat{p}$ $=\beta\Omega_{\tau}p\,,\quad\hat{m}=\alpha\sqrt{\beta\hat{\beta}}m_{\tau}\,,\quad\hat{q}=\alpha\beta q_{\tau}\,,\quad\hat{d}=\alpha\beta\left(\frac{d_{\tau}}{2}-\langle\tau^{2}\rangle_{\tau}\right)\,,$ (91)
$p=\big{\langle}s\langle\xi\rangle_{\xi|z,s,\hat{\xi}}\big{\rangle}_{z,s,\hat{\xi}}\qquad\qquad p_{\tau}=\big{\langle}s\langle\tau\rangle_{\tau|z,s,\hat{\tau}}\big{\rangle}_{z,s,\hat{\tau}}$ (92)
$q=\big{\langle}\langle\xi^{2}\rangle_{\xi|z,s,\hat{\xi}}\big{\rangle}_{z,s,\hat{\xi}}\qquad\qquad q_{\tau}=\big{\langle}\langle\tau\rangle_{\tau|z,s,\hat{\tau}}^{2}\big{\rangle}_{z,s,\hat{\tau}}$ (94)
@@
$P(\xi|z,s,\hat{\xi})=$ $P_{\xi}(\xi)\exp\left\{-\left(\frac{\hat{q}}{2}-\hat{d}\right)\xi^{2}+\xi\left(\hat{m}\hat{\xi}+\sqrt{\hat{q}}z+\hat{p}s\right)\right\}\,,$ (96)
$P(\tau|z,s,\hat{\tau})=$ $P_{\tau}(\tau)\exp\left\{-\left(\frac{\hat{q}_{\tau}}{2}-\hat{d_{\tau}}\right)\tau^{2}+\tau\left(\hat{m_{\tau}}\hat{\tau}+\sqrt{\hat{q_{\tau}}}z+\hat{p_{\tau}}s\right)\right\}\,,$ (97)
while the outher distributions $P_{\hat{\xi}}(\hat{\xi})$ and $P_{\hat{\tau}}(\hat{\tau})e^{\frac{\hat{\beta}}{2}(\hat{\tau})^{2}}$ depend from the choice of the planted configuration and $z$ is a Normal random variable.
@@ Appendix B RS equations in the interpolating case
So far, Eqs. (92-95) are general and valid for any arbitrary choice of the priors for all variables. In this section we derive explicitly the saddle point equations, replacing each of the random variables with their prior interpolation (7). The expressions of the distributions (96) and (97) can be recast as follows:
$Z_{\xi}$ ${}^{-1}\sum_{\epsilon}e^{-\frac{\xi^{2}}{2\gamma}+\xi\left(\phi\epsilon+h\right)}\,,$
$Z_{\tau}$ ${}^{-1}\sum_{\epsilon_{\tau}}e^{-\frac{\tau^{2}}{2\gamma_{\tau}}+\xi\left(\phi_{\tau}\epsilon_{\tau}+h_{\tau}\right)}\,,$
@@
$\gamma=\frac{\Omega_{\xi}}{1-\alpha\beta\Omega_{\xi}\left(d_{\tau}-q_{\tau}-\langle\tau^{2}\rangle_{\tau}\right)},\;\quad$ $\quad\gamma_{\tau}=\frac{\Omega_{\tau}}{1-\beta\Omega_{\tau}\left(d-q\right)}\,,$
$\phi=\frac{\sqrt{\delta}}{\Omega_{\xi}}\,,\quad$ $\quad\phi_{\tau}=\frac{\sqrt{\delta_{\tau}}}{\Omega_{\tau}}\,,$
$h=\hat{m}\hat{\xi}+\sqrt{\hat{q}}z+\hat{p}s\,,\quad$ $\quad h_{\tau}=\hat{m}_{\tau}\hat{\tau}+\sqrt{\hat{q}_{\tau}}z+\hat{p}_{\tau}s\,.$
The averages in the mean field equations (92-95) become:
$\langle\xi\rangle_{\xi|z,s,\hat{\xi}}=$ $\partial_{h}\ln Z_{\xi}=\frac{\sum_{\epsilon}e^{\frac{\gamma}{2}\left(\phi\epsilon+h\right)^{2}}\gamma\left(\phi\epsilon+h\right)}{\sum_{\epsilon}e^{\frac{\gamma}{2}\left(\phi\epsilon+h\right)^{2}}}=\gamma\,h+\gamma\phi\tanh{\left(\gamma\phi\,h\right)}$
@@
$\mathbb{E}_{s,\hat{\xi}}f(s,\hat{\xi})=\mathbb{E}_{g_{s,\hat{\xi}}}\mathbb{E}_{\epsilon_{s,\hat{\xi}}}f(\sqrt{\Omega_{s}}\,g_{s}+\sqrt{\delta_{s}}\epsilon_{s},\sqrt{\Omega_{\hat{\xi}}}\,g_{\hat{\xi}}+\sqrt{\delta_{\hat{\xi}}}\,\epsilon_{\hat{\xi}})\;.$
@@
$\mathbb{E}_{\hat{\tau}}\,e^{\frac{\hat{\beta}}{2}(\hat{\tau})^{2}}f(\hat{\tau})=\mathbb{E}_{\hat{g}_{\tau}}\mathbb{E}_{\hat{\epsilon}_{\tau}}\,e^{(\sqrt{\Omega_{\hat{\tau}}}\,g_{\hat{\tau}}+\sqrt{\delta_{\hat{\tau}}}\,\epsilon_{\hat{\tau}})^{2}}f(\sqrt{\Omega_{\hat{\tau}}}\,g_{\hat{\tau}}+\sqrt{\delta_{\hat{\tau}}}\,\epsilon_{\hat{\tau}})\,,$
@@
$\bar{m}_{\tau,*}$ $=\frac{1}{\mathbb{E}_{\hat{\tau}}\,e^{\frac{\hat{\beta}}{2}(\hat{\tau})^{2}}}\left\langle\mathbb{E}_{s}\,\mathbb{E}_{\hat{\tau}}\,e^{\frac{\hat{\beta}}{2}(\hat{\tau})^{2}}\epsilon_{\hat{\tau}}\tanh\left(\gamma_{\tau}\phi_{\tau}\left(\sqrt{\hat{\beta}\,\beta}\,m\hat{\tau}+\sqrt{\beta q}z+\beta\alpha\Omega_{\xi}p_{\tau}s\right)\right)\right\rangle_{z}$ (108)
$\bar{m}_{\tau,s}$ $=\frac{1}{\mathbb{E}_{\hat{\tau}}e^{\frac{\hat{\beta}}{2}(\hat{\tau})^{2}}}\left\langle\mathbb{E}_{s}\,\mathbb{E}_{\hat{\tau}}e^{\frac{\hat{\beta}}{2}(\hat{\tau})^{2}}\epsilon_{\tau,s}\tanh\left(\gamma_{\tau}\phi_{\tau}\left(\sqrt{\hat{\beta}\,\beta}\,m\hat{\tau}+\sqrt{\beta q}z+\beta_{2}\alpha\Omega_{\xi}p_{\tau}s\right)\right)\right\rangle_{z}$
$\bar{q}=$ $\frac{1}{\mathbb{E}_{\hat{\tau}}e^{\frac{\hat{\beta}}{2}(\hat{\tau})^{2}}}\left\langle\mathbb{E}_{s}\,\mathbb{E}_{\hat{\tau}}e^{\frac{\hat{\beta}}{2}(\hat{\tau})^{2}}\tanh^{2}\left(\gamma_{\tau}\phi_{\tau}\left(\sqrt{\hat{\beta}\,\beta}\,m\hat{\tau}+\sqrt{\beta q}z+\beta\alpha\Omega_{\xi}p_{\tau}s\right)\right)\right\rangle_{z}\,.$

### v2

@@ Appendix A Derivation of the RS equations
$\hat{f}^{RS}(\Lambda,\hat{\Lambda},\Lambda_{\tau},\hat{\Lambda}_{\tau})=\hat{f}^{RS}_{0}+n\hat{f}^{RS}_{1}+\mathcal{O}(n^{2})\,$,
$\hat{f}^{RS}_{0}$ $=\frac{\alpha\hat{\beta}}{2}d_{\hat{\tau}}d_{\hat{\xi}}+\hat{\beta}\operatorname{\mathrm{Extr}}_{\bm{p}_{T}}f_{T}(\bm{p}_{T})+\alpha\log\mathbb{E}_{\hat{\tau}}\exp\left(\hat{d}_{\hat{\tau}}(\hat{\tau})^{2}\right)+\log\mathbb{E}_{\hat{\xi}}\exp\left(\hat{d}_{\hat{\xi}}(\hat{\xi})^{2}\right)-\alpha\hat{d}_{\hat{\tau}}d_{\hat{\tau}}-\hat{d}_{\hat{\xi}}d_{\hat{\xi}}\;$,
$\hat{f}^{RS}_{1}$ $=\frac{1}{2}\hat{q}q+\alpha\frac{1}{2}\hat{q}_{\tau}q_{\tau}-\hat{m}m-\alpha\hat{m}_{\tau}m_{\tau}-\hat{d}d-\alpha\hat{d}_{\tau}d_{\tau}+\alpha^{2}\frac{\beta}{2}\Omega_{\xi}p_{\text{$\tau$}}^{2}+\frac{\beta}{2}\Omega_{\tau}p^{2}+$
$+\frac{\alpha}{\mathbb{E}_{\hat{\tau}}\exp\left(\hat{d}_{\hat{\tau}}\hat{\tau}^{2}\right)}\Bigg\langle\mathbb{E}_{\hat{\tau}}\exp\left(\hat{d}_{\hat{\tau}}\hat{\tau}^{2}\right)\mathbb{E}_{s}\log\mathbb{E}_{\tau}\exp\Bigg(-(\frac{\hat{q}_{\tau}}{2}-\hat{d}_{\tau})\tau^{2}+\tau(\hat{m}_{\tau}\hat{\tau}+\sqrt{\hat{q}_{\tau}}z+\hat{p}_{\tau}s)\Bigg)\Bigg\rangle_{z}+$
$+\frac{1}{\mathbb{E}_{\hat{\xi}}\exp\left(\hat{d}_{\hat{\xi}}\hat{\xi}^{2}\right)}\Bigg\langle\mathbb{E}_{\hat{\xi}}\exp\left(\hat{d}_{\hat{\xi}}\,\hat{\xi}^{2}\right)\mathbb{E}_{s}\log\mathbb{E}_{\xi}\exp\Bigg(-(\frac{\hat{q}}{2}-\hat{d})\xi^{2}+\xi(\hat{m}\hat{\xi}+\sqrt{\hat{q}}z+\hat{p}s)\Bigg)\Bigg\rangle_{z}+$
@@
The extremization of $\hat{f}^{RS}_{0}$ involves only the teacher order parameters. The extremizers should be s.t. $\operatorname{\mathrm{Extr}}\hat{f}^{RS}_{0}=0$ in order to avoid divergences when $n\to 0$. Indeed, remembering Eqs.(70), the parameters of order zero in $n$ satisfy the following
$\hat{d}_{\hat{\xi}}$ $=0\,,\qquad d_{\hat{\xi}}=1\,$, (90)
$\hat{d}_{\hat{\tau}}$ $=\frac{\hat{\beta}}{2}\,,\qquad d_{\hat{\tau}}=\langle\hat{\tau}^{2}\rangle_{\hat{\tau}}\,$,
which implies $\hat{f}^{RS}_{0}=0$ and ensure the student partition function is not diverging. Using Eqs.(79) and updating $\hat{f}_{1}^{RS}$ with (90) we obtain the values of $\Lambda$ and $\Lambda_{\tau}$ that extremize $\hat{f}_{1}^{RS}$
$\hat{p}_{\tau}$ $=\beta\alpha\Omega_{\xi}p_{\tau}\,,\qquad\hat{m}_{\tau}=\sqrt{\beta\hat{\beta}}m\,,\qquad\hat{q}_{\tau}=\beta q\,,\qquad\hat{d}_{\tau}=\beta\frac{d}{2}\,$, (91)
$\hat{p}$ $=\beta\Omega_{\tau}p\,,\qquad\hat{m}=\alpha\sqrt{\beta\hat{\beta}}m_{\tau}\,,\qquad\hat{q}=\alpha\beta q_{\tau}\,,\qquad\hat{d}=\alpha\beta\left(\frac{d_{\tau}}{2}-\langle\tau^{2}\rangle_{\tau}\right)\,$, (92)
$p=\big\langle s\langle\xi\rangle_{\xi|z,s,\hat{\xi}}\big\rangle_{z,s,\hat{\xi}}\hskip 18.49988pt\hskip 18.49988ptp_{\tau}=\big\langle s\langle\tau\rangle_{\tau|z,s,\hat{\tau}}\big\rangle_{z,s,\hat{\tau}}$ (93)
$m=\big\langle\hat{\xi}\langle\xi\rangle_{\xi|z,s,\hat{\xi}}\big\rangle_{z,s,\hat{\xi}}\hskip 18.49988pt\hskip 18.49988ptm_{\tau}=\big\langle\hat{\tau}\langle\tau\rangle_{\tau|z,s,\hat{\tau}}\big\rangle_{z,s,\hat{\tau}}$ (94)
$q=\big\langle\langle\xi^{2}\rangle_{\xi|z,s,\hat{\xi}}\big\rangle_{z,s,\hat{\xi}}\hskip 18.49988pt\hskip 18.49988ptq_{\tau}=\big\langle\langle\tau\rangle_{\tau|z,s,\hat{\tau}}^{2}\big\rangle_{z,s,\hat{\tau}}$ (95)
$d=\big\langle\langle\xi^{2}\rangle_{\xi|z,s,\hat{\xi}}\big\rangle_{z,s,\hat{\xi}}\hskip 18.49988pt\hskip 18.49988ptd_{\tau}=\big\langle\langle\tau^{2}\rangle_{\tau|z,s,\hat{\tau}}\big\rangle_{z,s,\hat{\tau}}\,$. (96)
@@
[…]
@@ Appendix B RS equations in the interpolating case
So far, Eqs. (93-96) are general and valid for any arbitrary choice of the priors for all variables. In this section we derive explicitly the saddle point equations, replacing each of the random variables with their prior interpolation (7). The expressions of the distributions (97) and (98) can be recast as follows:
$Z_{\xi}$ $^{-1}\sum_{\epsilon}\exp\left(-\frac{\xi^{2}}{2\gamma}+\xi\left(\phi\epsilon+h\right)\right)\,$,
$Z_{\tau}$ $^{-1}\sum_{\epsilon_{\tau}}\exp\left(-\frac{\tau^{2}}{2\gamma_{\tau}}+\xi\left(\phi_{\tau}\epsilon_{\tau}+h_{\tau}\right)\right)\,$,
@@
$\gamma=\frac{\Omega_{\xi}}{1-\alpha\beta\Omega_{\xi}\left(d_{\tau}-q_{\tau}-\langle\tau^{2}\rangle_{\tau}\right)},\;\qquad$ $\qquad\gamma_{\tau}=\frac{\Omega_{\tau}}{1-\beta\Omega_{\tau}\left(d-q\right)}\,$,
$\phi=\frac{\sqrt{\delta}}{\Omega_{\xi}}\,$$\qquad$ $\qquad\phi_{\tau}=\frac{\sqrt{\delta_{\tau}}}{\Omega_{\tau}}\,$,
$h=\hat{m}\hat{\xi}+\sqrt{\hat{q}}z+\hat{p}s\,$$\qquad$ $\qquad h_{\tau}=\hat{m}_{\tau}\hat{\tau}+\sqrt{\hat{q}_{\tau}}z+\hat{p}_{\tau}s\,.$
The averages in the mean field equations (93-96) become:
$\langle\xi\rangle_{\xi|z,s,\hat{\xi}}$= $\partial_{h}\ln Z_{\xi}=\frac{\sum_{\epsilon}\exp\left(\frac{\gamma}{2}\left(\phi\epsilon+h\right)^{2}\right)\gamma\left(\phi\epsilon+h\right)}{\sum_{\epsilon}\exp\left(\frac{\gamma}{2}\left(\phi\epsilon+h\right)^{2}\right)}=\gamma\,h+\gamma\phi\tanh{\left(\gamma\phi\,h\right)}$
@@
[…]
@@
[…]
@@
$\bar{m}_{\tau,*}$ $=\frac{1}{\mathbb{E}_{\hat{\tau}}\,\exp\left(\frac{\hat{\beta}}{2}(\hat{\tau})^{2}\right)}\left\langle\mathbb{E}_{s}\,\mathbb{E}_{\hat{\tau}}\,\exp\left(\frac{\hat{\beta}}{2}(\hat{\tau})^{2}\right)\epsilon_{\hat{\tau}}\tanh\left(\gamma_{\tau}\phi_{\tau}\left(\sqrt{\hat{\beta}\,\beta}\,m\hat{\tau}+\sqrt{\beta q}z+\beta\alpha\Omega_{\xi}p_{\tau}s\right)\right)\right\rangle_{z}$ (109)
$\bar{m}_{\tau,s}$ $=\frac{1}{\mathbb{E}_{\hat{\tau}}\exp\left(\frac{\hat{\beta}}{2}(\hat{\tau})^{2}\right)}\left\langle\mathbb{E}_{s}\,\mathbb{E}_{\hat{\tau}}\exp\left(\frac{\hat{\beta}}{2}(\hat{\tau})^{2}\right)\epsilon_{\tau,s}\tanh\left(\gamma_{\tau}\phi_{\tau}\left(\sqrt{\hat{\beta}\,\beta}\,m\hat{\tau}+\sqrt{\beta q}z+\beta_{2}\alpha\Omega_{\xi}p_{\tau}s\right)\right)\right\rangle_{z}$
$\bar{q}=$ $\frac{1}{\mathbb{E}_{\hat{\tau}}\exp\left(\frac{\hat{\beta}}{2}(\hat{\tau})^{2}\right)}\left\langle\mathbb{E}_{s}\,\mathbb{E}_{\hat{\tau}}\exp\left(\frac{\hat{\beta}}{2}(\hat{\tau})^{2}\right)\tanh^{2}\left(\gamma_{\tau}\phi_{\tau}\left(\sqrt{\hat{\beta}\,\beta}\,m\hat{\tau}+\sqrt{\beta q}z+\beta\alpha\Omega_{\xi}p_{\tau}s\right)\right)\right\rangle_{z}\,.$
Acknowledgement